OpenStudy (anonymous):

Barbara purchased a home entertainment center for $2,665 using an 18-month deferred payment plan with an interest rate of 24.36%. She did not make any payments during the deferment period. What will Barbara’s monthly payment be if she must pay off the home entertainment center within three years after the deferment period?

OpenStudy (anonymous):

the answer choices are 74.03 149.09 106.29 150.85

OpenStudy (amistre64):

interest build up over the defered period ... Bk^d is our starting balance. we then want to make payments for n periods which tend to build up into a geometric sum as: P(1-k^n)/(1-k) this gives us: Bn = Bk^dk^n - P(1-k^n)/(1-k) solving for P when Bn = 0 we get Bk^dk^n (1-k)/(1-k^n) = P

OpenStudy (amistre64):

so plugging on our values we get: 2665k^(18)k^(3*12)(1-k)/(1-k^(3*12)), k=1+.2436/12 http://www.wolframalpha.com/input/?i=2665k%5E%2818%29k%5E%283*12%29%281-k%29%2F%281-k%5E%283*12%29%29%2C+k%3D1%2B.2436%2F12